OpenStudy (cobaltthunder267):

Simple Algebra Question! Evaluate the exponential expression: 32^(-2/5) I've tried everything, flipping it, dropping the negative outside, using the fifth root, not using roots at all. It still says it's wrong :(

OpenStudy (cobaltthunder267):

The Answers that it's told me are wrong are: 4, -4, 1/2, -1/2, 1/5120, -1/5120

OpenStudy (loser66):

hehehe... I failed for this question also!! I got 1/4

OpenStudy (cobaltthunder267):

@Loser66 Yes! It worked. Would you mind explaining how you got it, please? =)

OpenStudy (loser66):

\(a^{-1}= \dfrac{1}{a}\) ok?

OpenStudy (loser66):

then \(32^{-2/5}=\dfrac{1}{32^{2/5}}\)

OpenStudy (loser66):

but \(32=2^5\), then \(\dfrac{1}{32^{2/5}}=\dfrac{1}{(2^5)^{2/5}}=\dfrac{1}{2^2}=\dfrac{1}{4}\)

OpenStudy (cobaltthunder267):

Oh! That makes tons more sense than the convoluted things I ended up trying. Thank you very much!